OpenStudy (anonymous):

(How Much Water must be added to 20 liters of 50% alcohol solution to Dilute it to 10%)

OpenStudy (amistre64):

a + b = t ; *-x xa + yb = zt -xa -xb =-xt xa+yb = zt ------------ b(y-x)= t(z-x) b@y = t(z-x)/(y-x) maybe

OpenStudy (amistre64):

water@0 = 20(.50 - .10)/(0-.10); chop off the negative and se if it works :)

OpenStudy (amistre64):

i get 80, maybe 80(0) + 20(.50) ?= 100(.10) 0 + 10 ?= 10 yep